Write the slope-intercept form of the equation of the line passing through the
point (-2,3) and paralley to the line y=-5x-4.

Answer by sarah_adam(201)

slope intercept form of equation is y = mx+c
where c = constant and m = slope
since the equation of the line is parallel to y = (-5)x+(-4)
so the equation will be y = = (-5)x+c
since the equation passes through point (-2,3)
-2 = (-5)(3)+c
-2 = -15+c
c = 15-2 = 13
therfore the required equation in slope intercep form is
y = (-5)x + 13